Assume an economy has no imports and no taxes. If the marginal propensity to save is 0.8, by how much would autonomous expenditure have to increase, to cause real gdp to rise by 100 billion


1
Expert's answer
2021-10-26T17:22:17-0400

If consumers spend 80 cents out of each dollar of disposable income, we can conclude that the government spending multiplier in a simple Keynesian model is 20.

The multiplier will be 1 / (1 - MPC) or 1 / MPS = 1 / 0.2 = 5.

Autonomous expenditure multiplier will be 5/4, that is, an increase of $ 400 billion in autonomous expenditure  will lead to an increase in GDP of $ 500 billion. 

To cause real GDP to rise by 100 billion, autonomous expenditure have to increase of 80 billion.